# On the polarimetric response of the Nançay Radio Telescope and its impact on precision pulsar timing

### v1

@@ 4.1 Polarimetric profiles
We tested the calibration procedure presented in the previous section by analyzing NUPPI data on a sample of MSPs, whose names and main properties are listed in Table 3. The MSPs in this list are the same as those also used as test cases in Guillemot et al. (2023), with two exceptions: we introduced PSR J01252327 (Morello et al., 2019), a highly-stable MSP observed with the NRT since mid-2017 with high cadence, and discarded PSR J1713+0747 (Foster et al., 1993), which underwent very significant and abrupt changes in its pulse profile on MJD 59321 (see e.g. Xu et al., 2021; Singha et al., 2021; Jennings et al., 2022), rendering profile and timing analyses unnecessarily complex for the work presented here. Most of the MSPs listed in Table 3 were included in the EPTA and IPTA second data releases (Perera et al., 2019; EPTA Collaboration et al., 2023a) for gravitational wave searches.
@@
We selected NUPPI observations of the MSPs at 1.4 GHz, and prepared two datasets for each MSP. The first dataset consisted in observations calibrated with the IFA method (i.e., with the SingleAxis method of pcm), that is, using calibration parameters obtained from the analysis of the preceding observations of the noise diode, and assuming ideal polarization feeds and noise diode signal. In the second dataset, observations were calibrated using parameters derived from the analysis presented in Sect. 3. In practice, the Gaussian processes and archetypes for all calibration parameters were used to predict calibration solutions at the epochs of the MSP observations. In the rest of this section we will refer to this second dataset as the “METM” dataset, given that the polarization calibration was carried out using the Measurement Equation Template Matching method of PSRCHIVE.
@@
In Fig. 10 we show frequency-integrated polarimetric pulse profiles for PSR J17302304, from the two datasets. All 1.4 GHz NUPPI observations of the pulsar are shown in the figure. Similar figures for the MSPs listed in Table 3 can be found in Appendix A (see Figs. 16 to 19). It is clear, from the figures, that the polarimetric profiles for the METM dataset (displayed in the right-hand panels) are significantly more homogeneous than in the other dataset. The calibration procedure described in Sect. 3, based on the METM method, thus appears to have successfully “propagated” the improved calibration to the entire NUPPI dataset.
@@
In addition to comparing polarimetric profiles we investigated the impact of the new calibration scheme on the timing of these MSPs. We followed the same procedure as described in Guillemot et al. (2023) for the extraction of TOAs from the two datasets on each pulsar: observations were integrated in time and frequency, to form eight frequency sub-bands of 64 MHz each. Then, template profiles for each of the MSPs were constructed by summing the total intensity profiles for the ten highest S/N observations, and smoothing the results. TOAs were finally extracted by comparing observations with the template profiles, using the Fourier-domain Monte Carlo method implemented in the “FDM” algorithm of the pat tool of PSRCHIVE. In Fig. 11 we show ratios of S/N and TOA uncertainty values for PSR J17302304. Figs. 20 to 22 in Appendix B show the same figures for the other MSPs. Finally, in Table 3 we quote the median S/N and TOA uncertainty ratios, and . In almost all tested MSPs, the new polarization calibration scheme increased the S/N values of individual NUPPI observations, leading to values generally larger than 1. The largest values of are found for PSRs J17302304 and J17441134, which are strongly polarized MSPs, thus particularly susceptible to changes in the polarization calibration. The TOA uncertainty ratios, on the other hand, have median values smaller than 1 in most cases, as a result from the generally larger S/N values in the METM dataset. It can be noted from Fig. 11 and Figs. 20 to 22 that the S/N and TOA uncertainty ratios all display sudden changes around MJD 58600, with, in general, larger S/N ratios and lower TOA uncertainty ratios after the event. This step coincides with the replacement of a pre-amplifier that occurred in 2019, which significantly modified the response of the NRT. The significantly different polarimetric response of the NRT is also visible in the left-hand panels of Fig. 10 and Figs. 16 to 19: most of these figures show two distinct modes for the polarimetric profiles (this is particularly visible in the linear polarization profiles). The two modes correspond to the data taken before and after the pre-amplifier change.
@@ 4.2 Timing
As mentioned in Sect. 4.1, TOA datasets were extracted using the FDM algorithm of pat, from NUPPI data calibrated with the IFA method on the one hand, and with the METM method on the other hand. The earlier datasets will be referred to as “IFA + FDM” datasets in the rest of the article, while the latter datasets will be referred to as “METM + FDM”. Additionally, we extracted TOAs from the same datasets using the MTM technique also implemented in pat, and using versions of the same template profiles that included all four Stokes parameters. The Stokes parameters are used by the MTM technique to model a transformation between the template profile and the considered observations, determining TOAs that are less likely to be biased by distorsions of the pulse profile caused by instrumental artefacts. Doing this, we obtained two other TOA datasets for each MSP. We will refer to these two other TOA datasets as “IFA + MTM” and “METM + MTM”.
Timing analyses were conducted using the TEMPO2 (Hobbs et al., 2006) and ENTERPRISE (Ellis et al., 2020) pulsar timing packages. ENTERPRISE was run using the PTMCMC Sampler (Ellis & van Haasteren, 2017), applied under a Bayesian approach to estimate posterior distributions for the sampled parameters. Topocentric TOAs were converted to Barycentric Coordinate Time (TCB) using the DE440 solar system ephemeris (Park et al., 2021). Best-fit timing models for the MSPs in Table 3 from the EPTA Data Release 2 (EPTA Collaboration et al., 2023a) were used as starting points for our timing analyses. For each pulsar and TOA dataset, we determined two white noise parameters: an error scale factor (EFAC, ) and an error added in quadrature (EQUAD, ), such that:
where denotes the uncertainty on the i-th TOA, and is the EFAC- and EQUAD-corrected TOA uncertainty. We also modeled an achromatic red noise component, in the form of a power law in the Power Spectral Density domain:
where represents the frequency of the signal, is the amplitude of the signal and is its spectral index. We sampled this spectrum using 30 Fourier coefficients.
@@
We note, first of all, that for all tested MSPs but one (PSR J17441134), the lowest values are found for the METM+MTM dataset. Interestingly, the lowest found in the case of PSR J17441134 is for METM+FDM, i.e., with the improved calibration and without using the MTM algorithm for extracting TOAs. The fact that, as was noted in Rogers et al. (2024), MTM is theoretically expected to perform worse than standard template matching for this highly polarized pulsar due to the multiple correlation between the phase shift and the parameters of the Jones transformation matrix may explain the lower value for this TOA dataset than for the others. We note also that in almost all cases, lower values are obtained for METM+FDM than for IFA+FDM datasets, and lower values are found for METM+MTM than for IFA+MTM datasets.
@@
Detailed comparisons of the , , , and parameters (displayed in Figs. 14 and 15), are made difficult by the fact that the best-fit values found for these parameters often have large uncertainties. Nevertheless, we find that in some cases (e.g., those of PSRs J01252327, J1022+1001, J17302304 or J1857+0943) significantly smaller levels of achromatic noise were found in the analysis of the METM+FDM, IFA+MTM and METM+MTM datasets than in that of the IFA+FDM datasets. The latter TOA datasets, extracted from NUPPI data calibrated with the IFA method and using standard template matching, therefore contain higher levels of red noise, likely caused by imperfect calibration, that the improved calibration scheme presented in Sect. 3 or the usage of the MTM method for determining TOAs attenuated at least partially.
@@ 5 Summary and conclusions
In this paper, we have presented the continuation of the work initiated in 2019, on the polarization calibration of 1.4 GHz pulsar observations recorded with the NUPPI backend of the NRT. Until November 2019, pulsar observations were calibrated using the IFA method. In Guillemot et al. (2023) we presented an improved polarization calibration method for NRT pulsar observations, based on the MEM analyses of observations of the bright, highly linearly polarized pulsar J07422822 in a special mode in which the feed horn rotates by 180∘ across the 1 h observation. Following the work presented in Guillemot et al. (2023), we conducted three series of observations of several bright pulsars covering wide declination ranges, to test whether the polarimetric response of the NRT depends on hour angle and declination. The analysis of the data taken as part of these three series of observations was conducted with a modified version of pcm, enabling the joint analysis of multiple pulsar observations at once, and allowing the differential gain, differential phase, feed orientation or feed ellipticity parameters to vary with hour angle and declination. By applying the best-fit calibration solutions as obtained from these analyses to standard mode observations of MSPs conducted at similar dates, we found that the simplest model, i.e., the model in which the above-mentioned calibration parameters do not vary with hour angle and declination, gave the best results. With the exception of the absolute gain, the instrumental response of the NRT therefore does not appear to depend on the observed direction.
@@
Since the first rotating horn observations of J07422822 were made in late-2019, and because the instrumental response of the NRT evolved with time, a different polarization calibration procedure needed to be developed for the calibration of pulsar observations conducted between the first NUPPI observations in August 2011 and late-2019. The solution found involved METM analyses of NUPPI observations of the bright pulsars J0953+0755 and J1136+1551, observed regularly between mid-2011 and 2024. The analysis of J1136+1551 observations enabled us to determine time segments during which the polarimetric response of the NRT appeared to be consistent. Then, based on the calibration parameters determined from METM analyses of J0953+0755 observations we determined “archetypes”, representing the frequency variations of the different calibration parameters over the individual time segments. We finally modeled the time variations of the scale and offset parameters needed to be applied to the parameter archetypes to fit individual METM results on J0953+0755, using Gaussian processes. These Gaussian processes and the parameter archetypes provide a way to account for the abrupt changes in the instrumental response of the NRT. In the future we will investigate the possibility of building calibration solutions based on METM results for more pulsars than J0953+0755 only. By providing denser measurements of the calibration parameters in time, such an analysis should further improve the efficiency of this calibration method.
@@
Overall, unlike Dey et al. (2024) but similarly to Rogers et al. (2024) we find that the MEM and METM polarization calibration methods produce significantly better results than the IFA method for NUPPI data. It should be noted that the NUPPI TOAs included in the EPTA Data Release 2 (EPTA Collaboration et al., 2023a), which will also be included in the future IPTA Data Release 3, were extracted using the MTM algorithm, as recommended by Rogers et al. (2024), from NRT observations calibrated with the IFA method for observations conducted before late-2019, and using the MEM method for observations made since. In future pulsar timing studies with the NRT as well as in future EPTA Data Releases we will adopt the new calibration method presented in this article, in addition to extracting TOAs with the MTM method.
The improved polarization calibration and more homogeneous profiles across the NUPPI dataset should, in principle, simplify polarization studies (e.g., RM measurements, analysis of polarimetric properties of pulsars, etc.) and significantly improve the results of wide-band template matching on NUPPI observations. In future work we will investigate the latter aspects, and will also look into applying a similar technique as presented in Sect. 3 for the polarization calibration of BON data. Between the first BON observations in 2004 and 2008, no calibration of the polarization data was performed. Then, starting from 2008, observations of the noise diode were conducted prior to each BON observation, so that BON observations could then get calibrated using the IFA approach. Past PTA analyses have shown that BON data display significant noise (see e.g. Lentati et al., 2016), that may well result from inadequate polarization calibration. We expect that an improved calibration model, similar to the one presented in Sect. 3, will diminish the noise present in the BON data, making pulsar timing analyses using BON data more sensitive and reliable.
@@ Acknowledgements.
- Voisin et al. (2024) Voisin, G., Cognard, I., Saillenfest, M., et al. 2024, arXiv e-prints, arXiv:2411.10066
@@
Appendix A Polarimetric profiles
In this section we present 1.4 GHz NUPPI polarimetric pulse profiles for the selection of MSPs considered in our calibration tests (see Sect.3). In these profiles, shown in Figs. 16 to 19, the total intensity , linear polarization and circular polarization are respectively shown as black, red and blue lines. Left-hand panels show the profiles as obtained after calibrating the NUPPI data using the IFA calibration method. Right-hand panels show the same observations calibrated using the method presented in Sect. 3, which is based on the Measurement Equation Template Matching (METM) method of PSRCHIVE (see van Straten 2013).
All NUPPI observations of the considered pulsars are shown in these figures, and were normalized to the maximum value of the total intensity. The individual 2048-bin profiles were plotted with a transparency factor proportional to , with the S/N of the i-th observation and the maximum S/N value, and such that a factor of 0 would correspond to a fully transparent (i.e., invisible) profile. The profiles shown in Figs. 16 to 19 were not corrected for Faraday rotation caused by magnetic fields along the line-of-sight to the pulsars, across the frequency bandwidth of 512 MHz.
Appendix B Comparison of S/N values and TOA uncertainties

### v2

@@ 4.1 Polarimetric profiles
We tested the calibration procedure presented in the previous section by analyzing NUPPI data on a sample of MSPs, whose names and main properties are listed in Table 3. The MSPs in this list are the same as those also used as test cases in Guillemot et al. (2023), with two exceptions: we introduced PSR J01252327 (Morello et al., 2019), a highly stable MSP observed with the NRT since mid-2017 with high cadence, and discarded PSR J1713+0747 (Foster et al., 1993), which underwent very significant and abrupt changes in its pulse profile on MJD 59321 (see, e.g., Xu et al., 2021; Singha et al., 2021; Jennings et al., 2022), rendering profile and timing analyses unnecessarily complex for the work presented here. Most of the MSPs listed in Table 3 were included in the EPTA and IPTA second data releases (Perera et al., 2019; EPTA Collaboration et al., 2023a) for gravitational wave searches.
@@
We selected NUPPI observations of the MSPs at 1.4 GHz, and prepared two datasets for each MSP. The first dataset consisted in observations calibrated with the IFA method (i.e., with the SingleAxis method of pcm), that is, using calibration parameters obtained from the analysis of the preceding observations of the noise diode, and assuming ideal polarization feeds and noise diode signal. In the second dataset, observations were calibrated using parameters derived from the analysis presented in Sect. 3. In practice, the Gaussian processes and archetypes for all calibration parameters were used to predict calibration solutions at the epochs of the MSP observations. In the rest of this section we refer to this second dataset as the “METM” dataset, given that the polarization calibration was carried out using the METM method of PSRCHIVE.
@@
In Fig. 10 we show frequency-integrated polarimetric pulse profiles for PSR J17302304, from the two datasets. All 1.4 GHz NUPPI observations of the pulsar are shown in the figure. Similar figures for the MSPs listed in Table 3 can be found in Appendix A (see Figs. 16 to 19). It is clear, from the figures, that the polarimetric profiles for the METM dataset (displayed in the right panels) are significantly more homogeneous than in the other dataset. The calibration procedure described in Sect. 3, based on the METM method, thus appears to have successfully propagated the improved calibration to the entire NUPPI dataset.
@@
In addition to comparing polarimetric profiles we investigated the impact of the new calibration scheme on the timing of these MSPs. We followed the same procedure as described in Guillemot et al. (2023) for the extraction of TOAs from the two datasets on each pulsar: observations were integrated in time and frequency, to form eight frequency sub-bands of 64 MHz each. Then, template profiles for each of the MSPs were constructed by summing the total intensity profiles for the ten highest S/N observations, and smoothing the results. TOAs were finally extracted by comparing observations with the template profiles, using the Fourier-domain Monte Carlo method implemented in the “FDM” algorithm of the pat tool of PSRCHIVE. In Fig. 11 we show ratios of S/N and TOA uncertainty values for PSR J17302304. Figs. 20 to 22 in Appendix B show the same figures for the other MSPs. Finally, in Table 3 we quote the median S/N and TOA uncertainty ratios, and . In almost all tested MSPs, the new polarization calibration scheme increased the S/N values of individual NUPPI observations, leading to values generally higher than 1. The highest values of are found for PSRs J17302304 and J17441134, which are strongly polarized MSPs, thus particularly susceptible to changes in the polarization calibration. The TOA uncertainty ratios, on the other hand, have median values smaller than 1 in most cases, as a result from the generally higher S/N values in the METM dataset. It can be noted from Fig. 11 and Figs. 20 to 22 that the S/N and TOA uncertainty ratios all display sudden changes around MJD 58600, with, in general, higher S/N ratios and lower TOA uncertainty ratios after the event. This step coincides with the replacement of a pre-amplifier that occurred in 2019, which significantly modified the response of the NRT. The significantly different polarimetric response of the NRT is also visible in the left panels of Fig. 10 and Figs. 16 to 19: most of these figures show two distinct modes for the polarimetric profiles (this is particularly visible in the linear polarization profiles). The two modes correspond to the data taken before and after the pre-amplifier change.
@@ 4.2 Timing
As mentioned in Sect. 4.1, TOA datasets were extracted using the FDM algorithm of pat, from NUPPI data calibrated with the IFA method on the one hand, and with the METM method on the other hand. The earlier datasets are referred to as “IFA + FDM” datasets in the rest of the article, while the latter datasets are referred to as “METM + FDM”. Additionally, we extracted TOAs from the same datasets using the MTM technique also implemented in pat, and using versions of the same template profiles that included all four Stokes parameters. The Stokes parameters are used by the MTM technique to model a transformation between the template profile and the considered observations, determining TOAs that are less likely to be biased by distorsions of the pulse profile caused by instrumental artifacts. Doing this, we obtained two other TOA datasets for each MSP. We refer to these two other TOA datasets as “IFA + MTM” and “METM + MTM”.
Timing analyses were conducted using the TEMPO2 (Hobbs et al., 2006) and ENTERPRISE (Ellis et al., 2020) pulsar timing packages. ENTERPRISE was run using the PTMCMC Sampler (Ellis & van Haasteren, 2017), applied under a Bayesian approach to estimate posterior distributions for the sampled parameters. Topocentric TOAs were converted to Barycentric Coordinate Time (TCB) using the DE440 solar system ephemeris (Park et al., 2021). Best-fit timing models for the MSPs in Table 3 from the EPTA Data Release 2 (EPTA Collaboration et al., 2023a) were used as starting points for our timing analyses. For each pulsar and TOA dataset, we determined two white noise parameters: an error scale factor (EFAC, ) and an error added in quadrature (EQUAD, ), such that
where denotes the uncertainty on the -th TOA, and is the EFAC- and EQUAD-corrected TOA uncertainty. We also modeled an achromatic red noise component, in the form of a power law in the power spectral density domain:
Here represents the frequency of the signal, is the amplitude of the signal, and is its spectral index. We sampled this spectrum using 30 Fourier coefficients.
@@
We note, first of all, that for all tested MSPs but one (PSR J17441134), the lowest values are found for the METM+MTM dataset. Interestingly, the lowest found in the case of PSR J17441134 is for METM+FDM, i.e., with the improved calibration and without using the MTM algorithm for extracting TOAs. The fact that, as was noted in Rogers et al. (2024), MTM is theoretically expected to perform worse than standard template matching for this highly polarized pulsar due to the multiple correlation between the phase shift and the parameters of the Jones transformation matrix may explain the lower value for this TOA dataset than for the others. We also note that in almost all cases, lower values are obtained for METM+FDM than for IFA+FDM datasets, and lower values are found for METM+MTM than for IFA+MTM datasets.
@@
Detailed comparisons of the , , , and parameters (displayed in Figs. 14 and 15), are made difficult by the fact that the best-fit values found for these parameters often have large uncertainties. Nevertheless, we find that in some cases (e.g., those of PSRs J01252327, J1022+1001, J17302304 or J1857+0943) significantly lower levels of achromatic noise were found in the analysis of the METM+FDM, IFA+MTM and METM+MTM datasets than in that of the IFA+FDM datasets. The latter TOA datasets, extracted from NUPPI data calibrated with the IFA method and using standard template matching, therefore contain higher levels of red noise, likely caused by imperfect calibration, that the improved calibration scheme presented in Sect. 3 or the usage of the MTM method for determining TOAs attenuated at least partially.
@@ 5 Summary and conclusions
In this paper, we have presented the continuation of the work initiated in 2019, on the polarization calibration of 1.4 GHz pulsar observations recorded with the NUPPI backend of the NRT. Until November 2019, pulsar observations were calibrated using the IFA method. In Guillemot et al. (2023) we presented an improved polarization calibration method for NRT pulsar observations, based on the MEM analyses of observations of the bright, highly linearly polarized pulsar J07422822 in a special mode in which the feed horn rotates by 180∘ across the 1 h observation. Following the work presented in Guillemot et al. (2023), we conducted three series of observations of several bright pulsars covering wide declination ranges, to test whether the polarimetric response of the NRT depends on hour angle and declination. The analysis of the data taken as part of these three series of observations was conducted with a modified version of pcm, enabling the joint analysis of multiple pulsar observations at once, and allowing the differential gain, differential phase, feed orientation, or feed ellipticity parameters to vary with hour angle and declination. By applying the best-fit calibration solutions as obtained from these analyses to standard mode observations of MSPs conducted at similar dates, we found that the simplest model, i.e., the model in which the above-mentioned calibration parameters do not vary with hour angle and declination, gave the best results. With the exception of the absolute gain, the instrumental response of the NRT therefore does not appear to depend on the observed direction.
@@
Since the first rotating horn observations of J07422822 were made in late 2019, and because the instrumental response of the NRT evolved with time, a different polarization calibration procedure needed to be developed for the calibration of pulsar observations conducted between the first NUPPI observations in August 2011 and late 2019. The solution found involved METM analyses of NUPPI observations of the bright pulsars J0953+0755 and J1136+1551, observed regularly between mid-2011 and 2024. The analysis of J1136+1551 observations enabled us to determine time segments during which the polarimetric response of the NRT appeared to be consistent. Then, based on the calibration parameters determined from METM analyses of J0953+0755 observations we determined “archetypes”, representing the frequency variations of the different calibration parameters over the individual time segments. We finally modeled the time variations of the scale and offset parameters needed to be applied to the parameter archetypes to fit individual METM results on J0953+0755, using Gaussian processes. These Gaussian processes and the parameter archetypes provide a way to account for the abrupt changes in the instrumental response of the NRT. In the future we will investigate the possibility of building calibration solutions based on METM results for more pulsars than J0953+0755 only. By providing denser measurements of the calibration parameters in time, such an analysis should further improve the efficiency of this calibration method.
@@
Overall, unlike Dey et al. (2024) but similarly to Rogers et al. (2024) we find that the MEM and METM polarization calibration methods produce significantly better results than the IFA method for NUPPI data. It should be noted that the NUPPI TOAs included in the EPTA Data Release 2 (EPTA Collaboration et al., 2023a), which will also be included in the future IPTA Data Release 3, were extracted using the MTM algorithm, as recommended by Rogers et al. (2024), from NRT observations calibrated with the IFA method for observations conducted before late 2019, and using the MEM method for observations made since. In future pulsar timing studies with the NRT as well as in future EPTA Data Releases we will adopt the new calibration method presented in this article, in addition to extracting TOAs with the MTM method.
The improved polarization calibration and more homogeneous profiles across the NUPPI dataset should, in principle, simplify polarization studies (e.g., RM measurements, analysis of polarimetric properties of pulsars) and significantly improve the results of wide-band template matching on NUPPI observations. In future work we will investigate the latter aspects, and will also look into applying a similar technique as presented in Sect. 3 for the polarization calibration of BON data. Between the first BON observations in 2004 and 2008, no calibration of the polarization data was performed. Then, starting from 2008, observations of the noise diode were conducted prior to each BON observation, so that BON observations could then get calibrated using the IFA approach. Past PTA analyses have shown that BON data display significant noise (see, e.g., Lentati et al., 2016), that may well result from inadequate polarization calibration. We expect that an improved calibration model, similar to the one presented in Sect. 3, will diminish the noise present in the BON data, making pulsar timing analyses using BON data more sensitive and reliable.
@@ Acknowledgements.
- Voisin et al. (2025) Voisin, G., Cognard, I., Saillenfest, M., et al. 2025, A&A, 693, A143
@@
Appendix A Polarimetric profiles
In this section we present 1.4 GHz NUPPI polarimetric pulse profiles for the selection of MSPs considered in our calibration tests (see Sect.3). In these profiles, shown in Figs. 16 to 19, the total intensity , linear polarization and circular polarization are respectively shown as black, red and blue lines. Left panels show the profiles as obtained after calibrating the NUPPI data using the IFA calibration method. Right panels show the same observations calibrated using the method presented in Sect. 3, which is based on the METM method of PSRCHIVE (see van Straten 2013).
All NUPPI observations of the considered pulsars are shown in these figures, and were normalized to the maximum value of the total intensity. The individual 2048-bin profiles were plotted with a transparency factor proportional to , with the S/N of the -th observation and the maximum S/N value, and such that a factor of 0 would correspond to a fully transparent (i.e., invisible) profile. The profiles shown in Figs. 16 to 19 were not corrected for Faraday rotation caused by magnetic fields along the line-of-sight to the pulsars, across the frequency bandwidth of 512 MHz.
Appendix B Comparison of S/N values and TOA uncertainties
Appendix C Results of the timing analysis
In Sect. 4.2 we investigated the influence of the new polarization calibration method on pulsar timing quality, by analyzing NUPPI timing data on a selection of millisecond pulsars. In Table LABEL:tab:timing we provide the full list of best-fit parameters found for each pulsar and TOA dataset. Full descriptions of the analysis and of datasets can be found in Sect. 4.2.